Suppose T is a normal operator on V. Suppose also that v, w "\in" V satisfy the equations ||v||=||w||=2, Tv=3v, Tw=4w. Show that ||T(v+w)=10


1
Expert's answer
2021-07-26T11:21:18-0400

given that ||v|| = 2 = ||w||

T(v) = 3v and T(w) = 4w

since v, w are eigenvectors, with distinct eigenvalue, therefore  v is orthogonal to w

"\implies" 3v and 4w are also orthogonal

"\implies""\langle" T(v), T(w) "\rangle" = "\langle" 3v, 4w "\rangle" = 0


now consider,

"\langle" T(v + w), T(v + w) "\rangle" = "\langle" T(v) + T(w)), T(v)+ T(w) "\rangle"

= "\langle" T(v), T(w) "\rangle" +"\langle" T(w), T(v) "\rangle" +"\langle" T(v), T(v) "\rangle" +"\langle" T(w), T(w) "\rangle"

= 0 + 0 + "\langle" 3v, 3v "\rangle" + "\langle" 4w + 4w "\rangle"

= 32"\langle" v, v "\rangle" + 42"\langle" w, w "\rangle"

= 32||v||2 + 42||w||2

= 9"\cdot" 22 + 16"\cdot" 22

= 36 + 64

= 100

"\implies" ||T(v + w)||2 = 100

"\implies" ||T(v + w)|| = 10